# TWISTED CONJUGACY IN DIHEDRAL ARTIN GROUPS I: TORUS KNOT GROUPS

### v2

@@ INTRODUCTION
Next we need to understand the complexity of finding a minimal length element in $\mathsf{CycGeo}_{\phi,\overline{(3)}}$, using Lemma 3.10. This only requires us to check for any $y^{\pm(k+1)}$ terms, and apply $\phi$-cyclic permutations and rewriting rules. The complexity of this step is bounded above by $2n$. Finally Step 4 runs in linear time from the discussion above, and so our adapted algorithm has complexity of $\mathcal{O}(n)$. ∎
@@
Let $u,v\in X^{\ast}$. Let $\psi\in\mathrm{Inn}(G(m))$, i.e. $\psi(w)=g^{-1}wg$ for some $g\in X^{\ast}$. By 2.8, we need to check two cases.
@@
    $u\sim_{\psi}v$. This case reduces to solving the conjugacy problem for $(gu,gv)$ (see 2.3). The conjugacy problem in dihedral Artin groups is solvable in linear time by 2.6.
@@
    $u\sim_{\psi\circ\phi}v$, where $\phi\in\mathrm{Out}(G(m))$ is of the form in Eq. 3. This case reduces to solving the $\mathrm{TCP}_{\phi}(G(m))$ with respect to $\phi$ (see 2.3). This is solvable in linear time by 3.20.
@@
When $m$ is odd, we recall that $|\mathrm{Out}(G(m))|=2$ by 2.8, so we can apply a similar technique as [9, Theorem 5.1] to prove Section 1.
@@
Let $\varphi_{1},\dots\varphi_{s}\in\mathrm{Aut}(G(m))$ be given, and consider
$A=\langle\varphi_{1},\dots\varphi_{s}\rangle\leq\mathrm{Aut}(G(m))$. For each $i=1,\dots,s$, compute $w_{i}\in G(m)$ such that
$\varphi_{i}=\gamma_{w_{i}}\phi_{i}$, where $\gamma_{w_{i}}\in\mathrm{Inn}(G(m))$ and $\phi_{i}\in\mathrm{Out}(G(m))$. Given two words $u,v\in X^{\ast}$, we want to decide whether $v\sim\varphi_{i}(u)$ for some $\varphi_{i}\in A$.
If $\phi_{i}$ is trivial for every $i$, then $A\leq\mathrm{Inn}(G(m))$, and so the set $\{\varphi_{i}(u)\mid\varphi_{i}\in A\}$ is a collection of conjugates of $u$. Here our problem reduces to deciding if $v$ is conjugate to $u$, which is decidable by 2.6. Otherwise, the set $\{\varphi_{i}(u)\mid\varphi_{i}\in A\}$ is a collection of conjugates of $u$ and $\phi(u)$, where $\phi(u)$ is defined in Eq. 3. Here, we need to decide whether $v$ is conjugate to either $u$ or $\phi(u)$, which is decidable by two applications of 2.6. ∎
Our final result is then immediate from Section 1, Section 1 and 4.1.
@@ Acknowledgements
The author would like to thank Laura Ciobanu and Fujii Michihiko for helpful discussions.

### v3

@@ INTRODUCTION
Next we need to understand the complexity of finding a minimal length element in $\mathsf{CycGeo}_{\phi,\overline{(3)}}$, using Lemma 3.10. This only requires us to check for any $y^{\pm(k+1)}$ terms, and apply $\phi$-cyclic permutations and rewriting rules. The complexity of this step is bounded above by $2n$.
Finally for Step 4, we use the fact that there exists a quadratic time algorithm to check if two words are cyclic permutations of each other via a two-way deterministic pushdown automaton (2DPDA) [1, Example 9.11], which can then be adapted to a linear time algorithm on a RAM machine [1, Theorem 9.10]. To check if $\overline{u}_{\min}\sim_{\phi}\overline{v}_{\min}$, we need to check that the $y$-terms are a $\phi$-cyclic permutation of each other, and that the differences in the $x$-terms differ by an even number. We therefore adapt our original algorithm on the 2DPDA as follows.
First, we adjust the input tape, so that we can determine if the $y$-terms are a $\phi$-cyclic permutation of each other. Then we add a counter which, when reading $x^{\pm 1}$ terms, keeps track of when they differ from the symbol on the pushdown list. The 2DPDA accepts a word if and only if the $y$-terms are a $\phi$-cyclic permutation of each other, and the counter is equal to 0(mod $0\ (\mathrm{mod}\ 2)$ after reading a word. This extra step of adding a counter to our 2DPDA can be achieved in constant time, and so Step 4 takes linear time overall on a RAM machine. ∎
ALGORITHM FOR THE TWISTED CONJUGACY PROBLEM
We now construct an algorithm to solve the $\mathrm{TCP}(G(m))$ for odd dihedral Artin groups. To achieve this, we first require a solution to another decision problem, known as the simultaneous conjugacy problem (SCP).
Definition 4.1.
Let $G=\langle X\rangle$. For fixed $k\in\mathbb{N}$, the $k$-simultaneous conjugacy problem takes as input two $k$-tuples $(y_{1},\dots,y_{k}),(z_{1},\dots,z_{k})$, where each $y_{i},z_{i}\in X^{\ast}$, and decides whether there exists an element $g\in G$ such that $g^{-1}y_{i}g=_{G}z_{i}$ for all $i=1,\dots,k$. We say this is an effective solution if such an algorithm also produces a conjugator $g\in G$.
This decision problem is also known as conjugacy for finite lists, and some results on this decision problem can be found in [15, 16]. Note there exists finitely presented groups with solvable conjugacy problem and unsolvable SCP [4]. The following result will be useful later.
Theorem 4.2.
The $k$-simultaneous conjugacy problem is solvable in dihedral Artin groups, for all $k\in\mathbb{N}$. Moreover, this solution is effective.
Proof.
Let $G(m)=\langle X\rangle$, where $X$ is the free product generating set, and let $(y_{1},\dots,y_{k})$, $(z_{1},\dots,z_{k})$ be $k$-tuples where each $y_{i},z_{i}\in X^{\ast}$. We first assume that each word $w\in X^{\ast}$ from these tuples is of the form $w=(w_{1},w_{2})$, where $w_{1}$ is a free product normal form modulo the centre, and $w_{2}$ is a power of the central generator.
Let $y_{i}=(y_{i_{1}},y_{i_{2}})$, $z_{i}=(z_{i_{1}},z_{i_{2}})$ for some $1\leq i\leq k$. By [13, Proposition 3.1], checking if $y_{i}$ and $z_{i}$ are conjugate is equivalent to checking that $y_{i_{2}}=z_{i_{2}}$, and that $y_{i_{1}}$ is a cyclic conjugate of $z_{i_{1}}$. In particular, the SCP for $G(m)$ is equivalent to solving the SCP for $\mathbb{Z}_{2}\ast\mathbb{Z}_{m}$ when $m$ is odd, or $\mathbb{Z}_{m}\ast\mathbb{Z}$ when $m$ is even. In both cases, these are hyperbolic groups, which have solvable SCP by [4, Theorem A]. Moreover, this solution is effective. ∎
The complexity for solving the SCP in hyperbolic groups was later improved in [5, Theorem 1]. We will use these complexity estimates to determine the overall complexity of the twisted conjugacy problem in odd dihedral Artin groups.
Proposition 4.3.
Let $G(m)=\langle X\rangle$ be an odd dihedral Artin group. Let $\varphi=\psi\circ\phi\in\mathrm{Aut}(G(m))$, where $\psi\in\mathrm{Inn}(G(m))$, $\phi\in\mathrm{Out}(G(m))$ such that for all $w\in G(m)$, $\psi(w)=g^{-1}wg$ for some $g\in G(m)$. Then there exists an algorithm which takes as input $\varphi\in\mathrm{Aut}(G(m))$, and determines the element $g\in G(m)$ defined by $\psi\in\mathrm{Inn}(G(m))$.
Moreover, the algorithm runs in time $\mathcal{O}(2\mu)$, where $\mu=4\cdot\mathrm{max}\{1,l(\varphi(x)),l(\varphi(y))\}$. In particular, the algorithm runs in linear time based on the generators $x,y\in X$ and the automorphism $\varphi\in\mathrm{Aut}(G(m))$ given.
Proof.
By 2.8, we have two cases to consider. To check if $\phi\in\mathrm{Out}(G(m))$ is trivial, we need to check if there exists $g\in G(m)$ such that $g^{-1}xg=_{G(m)}\varphi(x)$ and $g^{-1}yg=_{G(m)}\varphi(y)$. This is an example of the SCP in $G(m)$, for which we can determine the element $g\in G(m)$ by 4.2. Otherwise, suppose $\phi\in\mathrm{Out}(G(m))$ is of the form in Eq. 3. In this case, we need to check if there exists $g\in G(m)$ such that $g^{-1}\phi(x)g=g^{-1}x^{-1}g=_{G(m)}\varphi(x)$, and $g^{-1}\phi(y)g=g^{-1}y^{-1}g=_{G(m)}\varphi(y)$. Again the element $g\in G(m)$ can be determined using the SCP.
For complexity, the inputs of the SCP algorithm are $(x^{\varepsilon_{x}},y^{\varepsilon_{y}}),(\varphi(x),\varphi(y))$, where $\varepsilon_{x},\varepsilon_{y}=\pm 1$. By [5, Theorem 1], the algorithm has complexity as stated, noting that we have to account for all 4 cases for the different values of $\varepsilon_{x}$ and $\varepsilon_{y}$. ∎
@@
Let $u,v\in X^{\ast}$, and let $\varphi=\psi\circ\phi\in\mathrm{Aut}(G(m))$ as in 4.3 be our inputs. We want to decide whether there exists $w\in X^{\ast}$ such that $v=_{G(m)}\varphi(w)^{-1}uw$. We will use 2.3 to reduce the problem to twisted conjugacy with respect to outer automorphisms. Here $\psi\in\mathrm{Inn}(G(m))$, i.e. $\psi(w)=_{G(m)}g^{-1}wg$ for some $g\in G(m)$, which can be determined in linear time by 4.3. We need to check two cases:
@@
    $u\sim_{\psi}v$: Here $v=_{G(m)}\psi(w)^{-1}uw=g^{-1}w^{-1}guw$. Rearranging gives $gv=_{G(m)}w^{-1}(gu)w$, and so this case reduces to solving the conjugacy problem with respect to $(gu,gv)$. Since $g$ is known, this is solvable in linear time by [13, Proposition 3.1].
@@
    $u\sim_{\psi\circ\phi}v$, where $\phi\in\mathrm{Out}(G(m))$ is of the form in Eq. 3: Here $v=_{G}\psi(\phi(w))^{-1}uw=g^{-1}\phi(w)^{-1}guw$. Rearranging gives $gv=_{G}\phi(w)^{-1}(gu)w$, and so this case reduces to solving the twisted conjugacy problem $\mathrm{TCP}_{\phi}(G(m))$ with respect to $\phi\in\mathrm{Out}(G(m))$. This is solvable in linear time by 3.21.
@@
When $m$ is odd, we recall that $|\mathrm{Out}(G(m))|=2$ by 2.8, so we can apply a similar technique as [12, Theorem 5.1] to prove Section 1.
@@
Let $\varphi_{1},\dots\varphi_{s}\in\mathrm{Aut}(G(m))$ be given, and consider $A=\langle\varphi_{1},\dots\varphi_{s}\rangle\leq\mathrm{Aut}(G(m))$. For each $i=1,\dots,s$, compute $w_{i}\in G(m)$ such that $\varphi_{i}=\gamma_{w_{i}}\phi_{i}$, where $\gamma_{w_{i}}\in\mathrm{Inn}(G(m))$ and $\phi_{i}\in\mathrm{Out}(G(m))$ (see 4.3). Given two words $u,v\in X^{\ast}$, we want to decide whether $v\sim\varphi(u)$ for some $\varphi\in A$.
If $\phi_{i}$ is trivial, for every $i$, then $A\leq\mathrm{Inn}(G(m))$, and so the set $\{\varphi(u)\mid\varphi\in A\}$ is a collection of conjugates of $u$. Here our problem reduces to deciding if $v$ is conjugate to $u$, which is decidable by [13, Proposition 3.1]. Otherwise, the set $\{\varphi(u)\mid\varphi\in A\}$ is a collection of conjugates of $u$ and $\phi(u)$, where $\phi\in\mathrm{Out}(G(m))$ as in Eq. 3. If $m$ is odd, we need to decide whether $v$ is conjugate to either $u$ or $\phi(u)$, which is decidable by two applications of [13, Proposition 3.1]. ∎
Our final result is then immediate from Section 1, Section 1 and 5.1.
@@ Acknowledgements
The author would like to thank Laura Ciobanu and Fujii Michihiko for helpful discussions. They would also like to thank Andrew Duncan and Alessandro Sisto for their comments.